A small business borrows $40,000, part of the loan from the bank at the rate of 8% annual simple interest, and the rest from a credit card at the rate of 14% annual simple interest. The interest paid on both loans after one year totaled $3,800. How much was borrowed at each rate? Thank You
Answer by Greenfinch(383)

Amount from Bank is A so the amount on credit card is 40000 - A
Interest is 0.08A + 0.14(40000 - A) = 3800
0.08A + 5600 - 0.14A = 3800
- 0.06A = - 1800
A = 30 000
So the Bank loan is $30000 and the Credit Card is $10000
Interest is $2400 on the $30000 at 8% and £1400 on $10000 at 14%